Calculate the molar mass of the following substances.
Hydrochloric acid, HCl

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student


Hydrochloric acid, HCl = 1 × 1 + 1 × 35.5
= 1 + 35.5 = 36.5 g


Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Molar mass of HCl = Mass of H + Mass of Cl
= 1 + 35.5
= 36.5 g/mol
 
Hope it helps!
Thank you ✌️